# Consistency Models Made Easy

### v1

@@ 3.3 Tackling the "Curse of Consistency"
Training a model with this loss objective results in a denoising diffusion model [21]. This observation suggests a learning scheme that smoothly interpolates from DMs $\Delta t=t$ to CMs $\Delta t=\mathrm{d}t$ by gradually shrinking $\Delta t\to\mathrm{d}t$ during training. With this reasoning, diffusion pretraining can be considered as a special case of consistency training with a loose discretization of the consistency condition. In practice, we recommend initializing ECT with a pretrained diffusion model. Another benefit of this initialization is that during training, especially in the initial stages, it ensures good targets $f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r})$ in the loss objective. Finally, we note that our method works even if we train CMs from scratch with the objective in Eq. 12 (See the results of ablation study in Appendix B). However, initializing ECT with a pretrained diffusion model further improves efficiency, as noted in Sec. 4.
@@
We investigate the design principles of a continuous-time schedule whose "boundary" condition yields diffusion pretraining, i.e., constructing training pairs of $r=0$ for all $t$ at the beginning. Note that this is unlike the training schedule used in iCT (See Sec. 2 and Appendix A). We consider overlapped intervals for consistency training, which allows for factoring $p(t,r)=p(t)\,p(r|t)$ and continuous sampling of infinite $t$ from noise distribution $p(t)$, for instance, $\mathrm{LogNormal}(P_{\text{mean}},P_{\text{std}})$, and $r\sim p(r|t)$.
@@
Input: Dataset ${\mathcal{D}}$, a pretrained diffusion model $\theta_{\phi}$, mapping function $p(r\mid t,\mathrm{Iters})$, weighting function $w(t)$.
@@
As discussed in Sec. 2, iCT uses the Pseudo-Huber metric [7] to mitigate the perceptual bias caused by the LPIPS metric [76]. When taking a careful look as this metric, we reveal that this metric offers adaptive per-sample scaling of the gradients, which reduces the variance of gradients during training. Specifically, the differential of the Pseudo-Huber loss can be decomposed into two terms: an adaptive scaling factor $w(\Delta)$ and the differential of the squared $L_{2}$ loss. We discuss this in more detail in Appendix A. Therefore, we retain the squared $L_{2}$ metric used in DMs, and explore varying adaptive weighting terms.
@@
Weighting functions usually lead to a substantial difference in performance in DMs, and the same holds true for CMs. Substituting the mapping function $p(r|t)$ in Eq. 15 into the weighting function in Eq. 12, we get $w(t,r)=\nicefrac{{1}}{{(t-r)}}=q^{a}\cdot\nicefrac{{1}}{{tn(t)}}$, where $q^{a}$ is the scaling factor of the loss when $\Delta t\to 0$, and $\nicefrac{{1}}{{tn(t)}}$ is the weighting akin to DMs. This couples the weighting function with the mapping function. Instead, we consider decoupled weighting functions. Motivated by the adaptive scaling factor that appears in Pseudo-Huber loss (See Appendix A for more details), we rewrite the weighting function as
$w(t)=\bar{w}(t)\cdot w(\Delta)=\bar{w}(t)\cdot\frac{1}{\sqrt{\|\Delta\|_{2}^{2}+c^{2}}}$ (16)
where $\Delta=f({\mathbf{x}}_{t})-f({\mathbf{x}}_{r})$. We define $\bar{w}(t)$ as regular weighting, while $w(\Delta)$ as adaptive weighting. The adaptive weighting aids in improved training efficiency with the $L_{2}$ metric because when $\Delta\to 0$ (usually happens at $t\to 0$), this weighting $w(\Delta)$ can upscale gradients to avoid vanishing gradient in learning fine-grained features, while $c$ avoids potential numerical issues at $\Delta\approx 0$. We direct the reader to Tab. 2 and Tab. 3 in Appendix B for a detailed overview of various choices of $\bar{w}(t)$ and $w(\Delta)$ considered in this work. In general, we notice that CMs’ generative capability greatly benefits from weighting functions that control the variance of the gradients across different noise levels.
@@ 4 Experiments
We measure the efficiency and scalability of ECT on two datasets: CIFAR-10 [35] and ImageNet $64\times 64$ [10]. We evaluate the sample quality using Fréchet Inception Distance (FID) [18] and Fréchet Distance under the DINOv2 model [51] ($\text{FD}_{\text{DINOv2}}$ [68]) and measure sampling efficiency using the number of function evaluations (NFEs). We also indicate the relative training costs of each of these methods. The extensive implementation details can be found in Appendix C.
@@ 4.1 Comparison of Training Schemes
We compare CMs trained with ECT (denoted as ECM) against state-of-the-art diffusion models, diffusion models with advanced samplers, distillation methods such as consistency distillation (CD) [67], and improved Consistency Training (iCT) [64]. The key results are summarized in Fig. 2. We show the training FLOPs, inference cost, and generative performance of the four training schemes.
Figure 2: Comparison of training schemes for the diffusion-consistency family on CIFAR-10. Without relying on distillation from frozen diffusion teachers or extra adversarial supervision, ECT surpasses Consistency Distillation (CD) [67] and Consistency Models trained from scratch (iCT) [64] using $\nicefrac{{1}}{{32}}$ of the training cost. ECT also significantly reduces the inference cost to $\nicefrac{{1}}{{1000}}$ compared to Diffusion Pretraining (Score SDE/DMs) while maintaining comparable sample quality.
@@
Diffusion models.
We compare ECMs against Score SDE [66], EDM  [26], and EDM with DPM-Solver-v3 [79]. $2$-step ECM, which has been fine-tuned for only $100$k iterations, matches Score SDE-deep, with 2×2$\times$2 × model depth and $2000$ NFEs, in terms of FID. As noted in Fig. 2, ECM only requires $\nicefrac{{1}}{{1000}}$ of its inference cost and latency to achieve the same sample quality. $2$-step ECM fine-tuned for $100$k iterations outperforms EDM with advanced DPM-Solver-v3 (NFE=$10$).
@@
We compare ECT against Consistency Distillation (CD) [67], a SoTA approach that distills a pretrained DM into a CM. As shown in Tab. 1, ECM significantly outperforms CD on both CIFAR-10 and ImageNet 64$\times$64. We note that ECT is free from the errors of teacher DM and does not incur any additional cost of running teacher DM. $2$-step ECM outperforms $2$-step CD (with LPIPS [76]) in terms of FID ($2.20$ vs $2.93$) on CIFAR-10 while using around $\nicefrac{{1}}{{3}}$ of training/fine-tuning compute of CD.
@@
Improved Consistency Training (iCT) [64] is the SoTA recipe for training a consistency model from scratch without inferring the diffusion teacher. Compared to training from scratch, ECT rivals iCT-deep using $\nicefrac{{1}}{{32}}$ of the training compute and $\nicefrac{{1}}{{2}}$ of the model size as shown in Fig. 2 and Tab. 1.
@@
Diffusion Models
Score SDE [65] 2.38 2000
Score SDE-deep [65] 2.20 2000
EDM [26] 2.01 35
EDM (DPM-Solver-v3) [79] 2.51 10
Diffusion Distillation
PD [58] 8.34 1
GET [13] 5.49 1
Diff-Instruct [46] 4.53 1
TRACT [3] 3.32 2
CD (LPIPS) [67] 3.55 1
CD (LPIPS) [67] 2.93 2
Consistency Models
iCT [64] 2.83 1
iCT [64] 2.46 2
iCT-deep [64] 2.51 1
iCT-deep [64] 2.24 2
ECT
ECM (100k iters) 4.54 1
ECM (100k iters) 2.20 2
ECM (200k iters) 3.86 1
ECM (200k iters) 2.15 2
ECM (400k iters) 3.60 1
ECM (400k iters) 2.11 2
ImageNet 64$\times$64
@@
Diffusion Models
ADM [26] 2.07 250
EDM [26] 2.22 79
EDM2-XL [27] 1.33 63
Diffusion Distillation
BOOT [14] 16.3 1
DFNO (LPIPS) [78] 7.83 1
Diff-Instruct [46] 5.57 1
TRACT [3] 4.97 2
PD (LPIPS) [58] 5.74 2
CD (LPIPS) [67] 4.70 2
Consistency Models
iCT [64] 3.20 2
iCT-deep [64] 2.77 2
ECT
ECM-S (100k iters) 3.18 2
ECM-M (100k iters) 2.35 2
ECM-L (100k iters) 2.14 2
ECM-XL (100k iters) 1.96 2
ECM-S 4.05 1
ECM-S 2.79 2
ECM-XL 2.49 1
ECM-XL 1.67 2
ECT unlocks state-of-the-art few-step generative abilities through a simple yet principled approach. With a negligible tuning cost, ECT demonstrates strong results while benefiting from the scaling in training FLOPs to enhance its few-step generation capability.
@@ 4.2 Scaling Laws of ECT
We investigate the scaling behavior of ECT, focusing on three aspects: training compute, model size, and model FLOPs. When computing resources are not a bottleneck, ECT scales well and follows the classic power law.
@@
Initializing from the weights of EDM [26], we fine-tune ECMs across six compute scales on CIFAR-10 [35] and plot the trend of $\text{FD}_{\text{DINOv2}}$ against training compute in Fig. 4 (Left). The largest compute reaches 2×2$\times$2 × the diffusion pretraining budget. As we scale up the training budget, we observe a classic power-law decay in $\text{FD}_{\text{DINOv2}}$, indicating that increased computational investment in ECT leads to substantial improvements in generative performance. Intriguingly, the gap between 1-step and 2-step generation becomes narrower when scaling up training compute, even while using the same $\Delta t\to\mathrm{d}t$ schedule. We further fit the power-law $\text{FD}_{\text{DINOv2}}=K\cdot C^{\alpha}$, where $C$ is the normalized training FLOPs. The Pearson correlation coefficient between $\log(\text{Training Compute})$ and $\log(\text{FD}_{\text{DINOv2}})$ for 1-step and 2-step generation is $-0.9940$ and $-0.9996$, respectively, both with statistical significance ($p$-values of $5.385\times 10^{-5}$ and $2.798\times 10^{-7}$).
@@
Initializing from the weights of EDM2 [28], we train ECM-S/M/L/XL models, with parameters ranging from $280\mathrm{M}$ to $1.1\mathrm{B}$ and model FLOPs spanning from $102\mathrm{G}$ to $406\mathrm{G}$. As demonstrated in Fig. 4, both 1-step and 2-step generation capabilities exhibit log-linear scaling with respect to model FLOPs and parameters. This scaling behavior confirms that ECT effectively leverages increased model sizes and computational power to improve 1-step and 2-step generative performance. Notably, ECT achieves better 2-step generation performance than state-of-the-art CMs [64], while utilizing a remarkably modest budget of $12.8\mathrm{M}$ training images, which represents only $\mathbf{0.39\%}$ of the iCT [64] training budget, and $\mathbf{0.60\%}$ to $\mathbf{1.91\%}$ of the EDM2 [28] pretraining budget depending on the model sizes.
Figure 4: Left: Scaling up training compute yields the classic power-law between $\mathrm{FD}_{\text{DINOv2}}\downarrow$ and training compute. Right: Given the same batch size and iterations, scaling up model sizes and model FLOPs strongly correlates with $\mathrm{FID}\downarrow$ improvements on ImageNet $64\times 64$. The diameter is proportional to the model size.
@@ 5 Related Work
Consistency models [67, 64] are a new family of generative models designed for efficient generation with 1 or 2 model steps without the need for adversarial training. CMs do not rely on a pretrained diffusion model (DM) to generate training targets but instead leverage an unbiased score estimator. CMs have been extended to multi-step sampling [29], latent space models [45], ControlNet [71], and combined with extra adversarial loss [29, 33]. Despite their sampling efficiency, CMs are typically more challenging to train and require significantly more compute resources compared to their diffusion counterparts. Our work substantially improves the training efficiency of CMs, reducing the cost of future research and deployment on CMs.
Diffusion Distillation
@@
Drawing inspiration from knowledge distillation [19], diffusion distillation is the most widespread training-based approach to accelerate the diffusion sampling procedure. In diffusion distillation, a pretrained diffusion model (DM), which requires hundreds to thousands of model evaluations to generate samples, acts as a teacher. A student model is trained to match the teacher model’s sample quality, enabling it to generate high-quality samples in a few steps.
@@
There are two main lines of work in this area. The first category involves trajectory matching, where the student learns to match points on the teacher’s sampling trajectory. Methods in this category include offline distillation [44, 78, 13], which require an offline synthetic dataset generated by sampling from a pretrained DM to distill a teacher model into a few-step student model; progressive distillation [58, 48], and TRACT [3], which require multiple training passes or offline datasets to achieve the same goal; and BOOT [14], Consistency Distillation (CD)[67], and Imagine-Flash[32], which minimize the difference between the student predictions at carefully selected points on the sampling trajectory.
@@
CD is closely related to our method, as it leverages a teacher model to generate pairs of adjacent points and enforces the student predictions at these points to map to the initial data point. However, it limits the quality of consistency models to that of the pretrained diffusion model, and the LPIPS-based metric [76] used in CD loss introduces undesirable bias when evaluated on metrics such as FID.
@@
The second category minimizes the probabilistic divergence between data and model distributions, i.e., distribution matching [54, 70, 46, 74]. Methods such as DreamFusion [54] and ProfilicDreamer [70] are also used for 3D object generation. These methods [46, 74, 50, 60, 32, 72, 37] use score distillation or adversarial loss to distill an expensive teacher model into an efficient student model. However, they can be challenging to train in a stable manner due to the alternating updating schemes from either adversarial or score distillation.
@@
A drawback of training-based approaches is that they usually require extra training procedures after pretraining to distill an efficient student. For a detailed discussion on the recent progress of diffusion distillation, we refer to [12].
@@
Fast samplers are usually training-free and use advanced solvers to simulate the diffusion stochastic differential equation (SDE) or ordinary differential equation (ODE) to reduce the number of sampling steps. These methods reduce the discretization error during sampling by analytically solving a part of SDE or ODE [42, 43], by using exponential integrators and higher order polynomials for better approximation of the solution [75], using higher order numerical methods [26], using better approximation of noise levels during sampling [34], correcting predictions at each step of sampling [77] and ensuring that the solution of the ODE lies on a desired manifold [40]. Another orthogonal strategy is to achieve acceleration through parallelizing the sampling process [53, 62]. A drawback of these fast samplers is that the quality of samples drastically reduces as the number of sampling steps goes below a threshold such as $10$ steps.

### v2

@@ 3.3 Easy Consistency Tuning (ECT)
Training a model with this loss is identical to diffusion model/Score SDE (Ho et al., 2020, Song et al., 2021b). This observation suggests a learning scheme that smoothly interpolates from DMs $\Delta t=t$ to CMs $\Delta t=\mathrm{d}t$ by gradually shrinking $\Delta t\to\mathrm{d}t$ during training by gradually tightening the consistency condition. With this reasoning, diffusion pretraining can be considered as a special case of consistency training with a loose discretization of the consistency condition. Therefore, in practice, we start ECT with a pretrained diffusion model, resulting in a training scheme of pretraining+tuning. Another benefit of this initialization is that during training, especially in the initial stages, it ensures good targets $f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r})$ in the loss objective, avoiding trivial solutions.
We highlight two advantages of this learning scheme: 1. The pretraining+tuning scheme of ECT outperforms iCT’s training-from-scratch approach with lower overall computational cost (see Sec. 4.1). 2. Tuning serves as an efficient proxy for exploring the CM design space. Given a pretrained diffusion model, insights gained during tuning can be applied to improve the pretraining stage, resulting in a more refined overall CM training pipeline. Please refer to Appendix B for more details.
@@
Input: Dataset ${\mathcal{D}}$, a pretrained diffusion model $\phi$, mapping function $p(r\mid t,\mathrm{Iters})$, weighting function $w(t)$.
@@
We investigate the design principles of a continuous-time schedule whose "boundary" condition yields standard diffusion pretraining, i.e., constructing training pairs of $r=0$ for all $t$ at the beginning. Note that this is unlike the discrete-time schedule used in iCT (See Sec. 2 and Appendix A). We consider overlapped intervals for consistency models, which allows for factoring $p(t,r)=p(t)\,p(r|t)$ and continuous sampling of infinite $t$ from noise distribution $p(t)$, for instance, $\mathrm{LogNormal}(P_{\text{mean}},P_{\text{std}})$, and $r\sim p(r|t)$.
@@
iCT uses the Pseudo-Huber metric (Charbonnier et al., 1997) to mitigate the perceptual bias caused by the LPIPS metric (Zhang et al., 2018). When taking a careful look at this metric, we reveal that this metric offers adaptive per-sample scaling of the gradients, which reduces the variance of gradients during training. Specifically, the differential of the Pseudo-Huber loss can be decomposed into two terms: an adaptive scaling factor $w(\Delta)$ and the differential of the squared $L_{2}$ loss. We discuss this in more detail in Appendix A. Thus, we disentangle the weighting function from the loss metric as it provides greater flexibility. In our experiments, we retain the squared $L_{2}$ metric used in DMs and explore alternate choices of adaptive weighting terms that reduce the variance of gradients during training. Please refer to Appendix B for further discussions.
@@
Weighting functions usually lead to a substantial difference in performance in DMs, and the same holds true for CMs. Substituting the mapping function $p(r|t)$ in Eq. 15 into the weighting function in Eq. 12, we get $w(t,r)=\nicefrac{{1}}{{(t-r)}}=q^{a}\cdot\nicefrac{{1}}{{tn(t)}}$, where $q^{a}$ is the scaling factor to maintain the loss scales when $\Delta t\to 0$, and $\nicefrac{{1}}{{tn(t)}}$ is the weighting akin to DMs. This couples the weighting function with the mapping function $p(r|t)$. Instead, we consider decoupled weighting functions without relying on $p(r|t)$. Motivated by the adaptive scaling factor that appears in Pseudo-Huber loss (See Appendix A for more details), we rewrite the weighting function as
$w(t)=\bar{w}(t)\cdot w(\Delta)=\bar{w}(t)\cdot\frac{1}{(\|\Delta\|_{2}^{2}+c^{2})^{\nicefrac{{1}}{{2}}}},$ (16)
where $\Delta=f({\mathbf{x}}_{t})-f({\mathbf{x}}_{r})$. We define $\bar{w}(t)$ as timestep weighting, while $w(\Delta)$ as adaptive weighting. The adaptive weighting improves training efficiency with the $L_{2}$ metric because, as $\Delta\to 0$ (usually happens when $t\to 0$), this weighting $w(\Delta)$ upscales gradients to avoid vanishing gradient during learning of fine-grained features, while $c$ mitigates potential numerical issues when $\Delta\approx 0$. We direct the reader to Tab. 2 in Appendix B for a detailed overview of various choices of $\bar{w}(t)$ and $w(\Delta)$ considered in this work. In general, we notice that CMs’ generative capability greatly benefits from weighting functions that control the variance of the gradients across different noise levels. We provide a brief summary of the core steps of ECT in Alg. 1.
@@ 4 Experiments
This section compares different learning schemes and investigates scaling laws of ECT, while more experiments on design choices and scaling are shown in Appendix B. We evaluate the efficiency and scalability of ECT on two datasets: CIFAR-10 (Krizhevsky, 2009) and ImageNet $64\times 64$ (Deng et al., 2009). We measure the sample quality using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Distance under the DINOv2 model (Oquab et al., 2023) ($\text{FD}_{\text{DINOv2}}$(Stein et al., 2024) and sampling efficiency using the number of function evaluations (NFEs). We also indicate the relative training costs of each of these methods. Implementation details can be found in Appendix D.
@@ 4.1 Comparison of Training Schemes
We compare CMs trained with ECT (denoted as ECM) against state-of-the-art diffusion models, advanced samplers for diffusion models, distillation methods such as consistency distillation (CD) (Song et al., 2023), and improved Consistency Training (iCT) (Song and Dhariwal, 2023). The key results are summarized in Fig. 2. We show the training FLOPs, inference cost, and generative performance of the four training schemes.
Figure 2: Comparison of training schemes for the diffusion-consistency family on CIFAR-10. Without relying on distillation from frozen diffusion teachers or extra adversarial supervision, ECT surpasses Consistency Distillation (CD) (Song et al., 2023) and Consistency Models trained from scratch (iCT) (Song and Dhariwal, 2023) using $\nicefrac{{1}}{{4}}$ of the total training cost. ECT also significantly reduces the inference cost to $\nicefrac{{1}}{{1000}}$ compared to Diffusion Pretraining (Score SDE/DMs) while maintaining comparable sample quality.
@@
Score SDE/Diffusion Models.
We compare ECMs against Score SDE (Song et al., 2021b), EDM  (Karras et al., 2022), and EDM with DPM-Solver-v3 (Zheng et al., 2024). $2$-step ECM, which has been fine-tuned for only $100$k iterations, matches Score SDE-deep, with 2×2$\times$2 × model depth and $2000$ NFEs, in terms of FID. As noted in Fig. 2, ECM only requires $\nicefrac{{1}}{{1000}}$ of its inference cost and latency to achieve the same sample quality. $2$-step ECM fine-tuned for $100$k iterations outperforms EDM with advanced DPM-Solver-v3 (NFE=$10$).
@@
We compare ECT against Consistency Distillation (CD) (Song et al., 2023), a SoTA approach that distills a pretrained DM into a CM. As shown in Tab. 1, ECM significantly outperforms CD on both CIFAR-10 and ImageNet 64$\times$64. We note that ECT is free from the errors of teacher DM and does not incur any additional cost of running teacher DM. $2$-step ECM outperforms $2$-step CD (with LPIPS (Zhang et al., 2018)) in terms of FID ($2.20$ vs $2.93$) on CIFAR-10 while using around $\nicefrac{{1}}{{3}}$ of training compute of CD.
@@
Improved Consistency Training (iCT) (Song and Dhariwal, 2023) is the SoTA recipe for training a consistency model from scratch without inferring the diffusion teacher. Compared to training from scratch, ECT rivals iCT-deep using $\nicefrac{{1}}{{4}}$ of the overall training compute ($\nicefrac{{1}}{{8}}$ in the tuning stage) and $\nicefrac{{1}}{{2}}$ of the model size as shown in Fig. 2 and Tab. 1.
@@ 4.2 Scaling Laws of ECT
We leverage efficiency of ECT to examine the scaling behavior of CMs, including training compute, model size, and model FLOPs. We find that when computational resources are not a bottleneck, ECT scales well and follows the classic power law.
@@
Initializing from the weights of EDM (Karras et al., 2022), we fine-tune ECMs across six compute scales on CIFAR-10 (Krizhevsky, 2009) and plot the trend of $\text{FD}_{\text{DINOv2}}$ against training compute in Fig. 4 (Left). The largest compute reaches 2×2$\times$2 × the diffusion pretraining budget. As we scale up the training budget, we observe a classic power-law decay in $\text{FD}_{\text{DINOv2}}$, indicating that increased computational investment in ECT leads to substantial improvements in generative performance. Intriguingly, the gap between 1-step and 2-step generation becomes narrower when scaling up training compute, even while using the same $\Delta t\to\mathrm{d}t$ schedule. We further fit the power-law $\text{FD}_{\text{DINOv2}}=K\cdot C^{\alpha}$, where $C$ is the normalized training FLOPs. The Pearson correlation coefficient between $\log(\text{Training Compute})$ and $\log(\text{FD}_{\text{DINOv2}})$ for 1-step and 2-step generation is $-0.9940$ and $-0.9996$, respectively, both with statistical significance ($p$-values $<10^{-4}$).
@@
Initialized from EDM2 (Karras et al., 2024) pretraining, we train ECM-S/M/L/XL models with parameters from $280\mathrm{M}$ to $1.1\mathrm{B}$ and model FLOPs from $102\mathrm{G}$ to $406\mathrm{G}$. As demonstrated in Fig. 4, both 1-step and 2-step generation capabilities exhibit log-linear scaling for model FLOPs and parameters. This scaling behavior confirms that ECT effectively leverages increased model sizes and computational power to improve 1-step and 2-step generative capabilities.
[…]
Inference.
Our scaling study also indicates a sweet spot for the inference of CMs. On both CIFAR-10 and ImageNet 64$\times$64, there are 2-step inferences of smaller models surpassing 1-step inferences of larger models, e.g., $498\mathrm{M}$ ECM-M against $1.1\mathrm{B}$ ECM-XL. This calls for further studies of inference-optimal scaling and test-time compute scaling for visual generation.
Figure 4: (Left): Scaling up training compute yields the classic power-law between $\mathrm{FD}_{\text{DINOv2}}\downarrow$ and training compute, with $K=263$, $\alpha=-0.060$ for 1-step inference, and $K=164$, $\alpha=-0.028$ for 2-step inference, (Right): Given the same batch size and iterations, scaling up model sizes and model FLOPs strongly correlates with $\mathrm{FID}\downarrow$ improvements on ImageNet $64\times 64$. The diameter is proportional to the model size.
@@ 5 Related Work
Consistency models (Song et al., 2023, Song and Dhariwal, 2023) are a new family of generative models designed for efficient generation with few model steps without the need for adversarial training. CMs do not rely on a pretrained diffusion model (DM) to generate training targets but instead leverage an unbiased score estimator. Training CMs have been extended to multi-step sampling (Kim et al., 2024, Wang et al., 2024b, Heek et al., 2024), latent space models (Luo et al., 2023), ControlNet (Xiao et al., 2023), video (Wang et al., 2024c), and combined with an additional adversarial loss (Kim et al., 2024, Kong et al., 2023). Despite their sampling efficiency, CMs are typically more challenging to train and require significantly more compute resources than their diffusion counterparts. Our work substantially improves the training efficiency of CMs, reducing the cost of future research and deployment on CMs. Parallel to our work, Heek et al. (2024), Wang et al. (2024b) relax the single-step constraint of CMs and generalize them to multi-step variants, drawing inspiration from TRACT (Berthelot et al., 2023). Under the piecewise linear design, increasing the number of segments will also lead to diffusion pretraining, showing a different approach to connect CMs to the standard diffusion models.
Diffusion Distillation.
@@
Drawing inspiration from knowledge distillation (Hinton et al., 2015), distillation is the most widespread training-based approach to accelerate the diffusion sampling procedure. In diffusion distillation, a pretrained diffusion model (DM), which requires hundreds to thousands of model evaluations to generate samples, acts as a teacher. A student model is trained to match the teacher model’s sample quality, enabling it to generate high-quality samples in a few steps.
@@
There are two main lines of work in this area. The first category involves trajectory matching, where the student learns to match points on the teacher’s sampling trajectory. Methods in this category include offline distillation (Luhman and Luhman, 2021, Zheng et al., 2023, Geng et al., 2024), which require an offline synthetic dataset generated by sampling from a pretrained DM to distill a teacher model into a few-step student model; progressive distillation (Salimans and Ho, 2022, Meng et al., 2023), and TRACT (Berthelot et al., 2023), which require multiple training passes or offline datasets to achieve the same goal; and BOOT (Gu et al., 2023), Consistency Distillation (CD)(Song et al., 2023), and Imagine-Flash (Kohler et al., 2024), which minimize the difference between the student predictions at carefully selected points on the sampling trajectory.
@@
CD is closely related to our method, as it leverages a teacher model to generate pairs of adjacent points and enforces the student predictions at these points to map to the initial data point. However, it employs a fixed schedule derived from a specific sampler, which may introduce non-negligible discretization errors in approximating the consistency condition. It also limits the quality of consistency models to that of the pretrained diffusion model.
@@
The second category minimizes the probabilistic divergence between data and model distributions, i.e., distribution matching (Poole et al., 2022, Wang et al., 2024a, Luo et al., 2024, Yin et al., 2023, Zhou et al., 2024a). These methods (Luo et al., 2024, Yin et al., 2023, Nguyen and Tran, 2023, Sauer et al., 2023, Kohler et al., 2024, Xu et al., 2023a, Lin et al., 2024, Zhou et al., 2024b) use score distillation or adversarial loss, to distill an expensive teacher model into an efficient student model. However, they can be challenging to train in a stable manner due to the alternating updating schemes from either adversarial or score distillation. Some of these methods such as DreamFusion (Poole et al., 2022) and ProlificDreamer (Wang et al., 2024a) are used for 3D object generation.
@@
A drawback of training-based approaches is that they need additional training procedures after pretraining to distill an efficient student, which can be computationally intensive. For a detailed discussion on the recent progress of diffusion distillation, we direct the readers to Dieleman (2024).
@@
Fast samplers are usually training-free and use advanced solvers to simulate the diffusion stochastic differential equation (SDE) or ordinary differential equation (ODE) to reduce the number of sampling steps. These methods reduce the discretization error during sampling by analytically solving a part of SDE or ODE (Lu et al., 2022a; b, Xue et al., 2024), by using exponential integrators and higher order polynomials for better approximation of the solution (Zhang and Chen, 2022), using higher order numerical methods (Karras et al., 2022), using better approximation of noise levels during sampling (Kong and Ping, 2021), correcting predictions at each step of sampling (Zhao et al., 2024) and ensuring that the solution of the ODE lies on a desired manifold (Liu et al., 2022a). Another orthogonal strategy is the parallel sampling process (Pokle et al., 2022, Shih et al., 2024), solving fixed points of the entire ODE/SDE trajectories. A drawback of these fast samplers is that the quality of samples drastically reduces as the number of sampling steps falls below a threshold such as 10 steps.
6 Conclusion
We propose Easy Consistency Tuning (ECT), a simple yet efficient scheme for training consistency models. The resulting models, ECMs, unlock state-of-the-art few-step generative capabilities at a minimal tuning cost and are able to benefit from scaling. We have made our code available to ease future prototyping, studying, and deploying consistency models within the community.